OpenStudy (anonymous):

Simplify completely 18 s t to the fourth power over 52 s cubed t times 16 s cubed over 9 s squared t??? help please :x

OpenStudy (goldphenoix):

Can you write it numerical?

OpenStudy (anonymous):

18st^4 16s^3 ------- *-------- 52s^3t 9s^2t

OpenStudy (goldphenoix):

\[\large \large \frac{ 18st^4 }{ 52s^3t } \times \frac{ 16s^3 }{ 9s^2t }\] ^It looks like this?

OpenStudy (anonymous):

yes ^.^

OpenStudy (goldphenoix):

Alright. So what do you think you want to do first?

OpenStudy (anonymous):

simplify...i have no idea:/

OpenStudy (goldphenoix):

Do you know how to cross multiply?

OpenStudy (anonymous):

yes

OpenStudy (goldphenoix):

Alright, so what is 62s^3t * 16s^3 ?

OpenStudy (anonymous):

well i got 992s^3t...doubt thats right lol

OpenStudy (goldphenoix):

Let take it step-by-step. :) What is 62 * 16? It is 992. But what is s^3 * s^3? Do you remember the Laws Of Exponents?

OpenStudy (anonymous):

its suppose to be 52 not 62 >.< btw ..and no

OpenStudy (anonymous):

(52*16=832)

OpenStudy (goldphenoix):

Opps my bad.

OpenStudy (goldphenoix):

Alright, the Laws Of Exponents say "To multiply powers of the same base, add the exponents" In this case, 1 is the base. Do you understand? Btw, you should remember this, it is very important! So what is s^3 * s^3?

OpenStudy (anonymous):

its 1?

OpenStudy (goldphenoix):

The base is 1. So s^3 * s^3 can be written as 1s^3 * 1s^3. So what would it give us? It's not 1.

OpenStudy (anonymous):

2s^3?

OpenStudy (goldphenoix):

Well. According to the Laws Of Exponents, you add the exponents. So we have 3 and 3. What is 3+3?

OpenStudy (anonymous):

6

OpenStudy (goldphenoix):

Alright. So s^3 * s^3 = s^6. Make sense so far?

OpenStudy (anonymous):

ohhhh yes makes sense lol

OpenStudy (goldphenoix):

So far, we have 832s^6. But what about the "t"? Well, we just leave it as t because you can't multiply t by anything. So we should get 832s^6t. Now what's 18st^4 * 9s^2t ?

OpenStudy (goldphenoix):

18 * 9 does equal to 162. But I don't understand the ^10 part. So we have to do st^4 * s^2t. What does s * s^2 give you?

OpenStudy (anonymous):

1^6

OpenStudy (goldphenoix):

Not quite. The base is 1. So another way of writing it is 1s * 1s^2, They both have the same base. So according to the Laws of Exponents, we add the exponents. So what do we get?

OpenStudy (anonymous):

wouldnt you add 2 and 4

OpenStudy (goldphenoix):

What do you mean? I'm talking about "s * s^2"

OpenStudy (anonymous):

2^2

OpenStudy (goldphenoix):

Here's 2 big tips: When you're multiplying a coefficient by a coefficient (A number times a variable. Ex: 1x, 10y, 205a) there should always be a variable. So what happened to the variable s? o_O And 2nd step, when you see a number/variable without an exponent, ex: s, 10, 100, 1000, a, d... there's always an invisible 1. So s can be s^1. 10 can be 10^1. and so on. So s is s^1. s^1 * s^2 is?

OpenStudy (anonymous):

1s^2

OpenStudy (goldphenoix):

What's 1 + 2?

OpenStudy (anonymous):

3

OpenStudy (goldphenoix):

So s^1 * s^2 = s^3. :) Now we are left with t^4 and t. What does that give us?

OpenStudy (anonymous):

1^4

OpenStudy (goldphenoix):

Remember what I said before? What exponent does t has?

OpenStudy (anonymous):

2^4

OpenStudy (goldphenoix):

Listen to what I said. :) What exponent does t has? Does it have ^1, ^2, ^1000, ^-1?

OpenStudy (anonymous):

ehhhhhhhh idk i give up lol

OpenStudy (goldphenoix):

No no! Don't give up! Just read what I said before. :) The tip I gave you. It can certainly help you. :)

OpenStudy (anonymous):

invisible 1 lol

OpenStudy (goldphenoix):

Alright!!! Good job. So t is t^1. What is t^1 * t^4?

OpenStudy (anonymous):

t^5

OpenStudy (goldphenoix):

Good job!!! So now our if you combine them, it looks like: 162s^3t^5 = 832s^6t

OpenStudy (anonymous):

thats it?...

OpenStudy (goldphenoix):

\[\large \large 162s^3t^5 = 832s^6t\] (Just to make it look easier to see) Are we suppose to find a variable, or just simplify in general?

OpenStudy (anonymous):

simplify completely

OpenStudy (goldphenoix):

Hmm. I thought we were suppose to find the variable. But if it's just simplify completely, I guess this is simplify completely. Let me ask @ganeshie8

OpenStudy (anonymous):

okay :)

OpenStudy (goldphenoix):

Hmm. I think it means simplify as in simplify the fraction. o_o So let's do that, it's much easier. Alright?

OpenStudy (anonymous):

answer choices: a)8t^2/13s b)8t/13s^2 c)4t^2/13s d)4t/39s^2

OpenStudy (goldphenoix):

Lolol. You should've told me that. :| So we have to simplify the fraction first. Okay?

OpenStudy (goldphenoix):

What does 18/52 simplify to?

OpenStudy (anonymous):

9/26

OpenStudy (goldphenoix):

Now remember what he said before, right? When you multiply 2 variable with an exponent, you add the exponent, right? Well, for division, instead of adding, you subtract the exponent.

OpenStudy (goldphenoix):

s^4/s^3 *

OpenStudy (anonymous):

s^1

OpenStudy (goldphenoix):

Opps. My bad. Yes s^4/s^3 does give you s. I meant to say, what is s/s^3?

OpenStudy (anonymous):

-2?

OpenStudy (goldphenoix):

Wow! Yes! Really good.

OpenStudy (goldphenoix):

What is t^4/t ?

OpenStudy (anonymous):

t^3

OpenStudy (goldphenoix):

Alright. So: \[\large \large \frac{ 18st^4 }{ 52s^3t } = \frac{ 9s^-2t^3 }{ 26 }\] What about the other fraction. What does 16/9 simplify to?

OpenStudy (goldphenoix):

I just realize, I've been helping you for an hour. :O A lot of time-consuming. :)

OpenStudy (anonymous):

thats why i hate math it a waste of time :) and doesnt it stay the same?

OpenStudy (goldphenoix):

Which one? 16/9?

OpenStudy (anonymous):

yeah

OpenStudy (goldphenoix):

In improper fraction form, yes. But in mixed number form, what would it be?

OpenStudy (anonymous):

i dont know :x

OpenStudy (goldphenoix):

Alright. Let skip that. What does s^3/s^2?

OpenStudy (anonymous):

s^1

OpenStudy (goldphenoix):

Correct! Or just s. So: \[\large \large \frac{ 16s^3 }{ 9s^2t } = \frac{ 16s }{ 9t }\]

OpenStudy (goldphenoix):

Make sense so far?

OpenStudy (anonymous):

yes :)

OpenStudy (goldphenoix):

Now: \[\large \large \frac{ 9s^-2t^3 }{ 26 } * \frac{ 16s }{ 9t }\]

OpenStudy (goldphenoix):

Wait a few min. I have to do something.

OpenStudy (anonymous):

okay :)

OpenStudy (goldphenoix):

I have to go. But the answer is 8t^2/13s

OpenStudy (anonymous):

thank you so much for your help :)

OpenStudy (anonymous):

This is correct! The answer is A) 8t^2/13s